ALGEBRAIC ANALYSIS OF BENT-FROM-LINEAR TRANSITION INTENSITIES: THE EMISSION SPEC-TRUM OF METHINOPHOSPHIDE (HCP)

\underline{F. PÉREZ-BERNAL} , Departamento de Física Aplicada, Facultad de Ciencias Experimentales, Avda. de las FF.AA. s/n, Universidad de Huelva, 21071 Huelva, Spain; F. IACHELLO and P. H. VAC-CARO, Departments of Physics and Chemistry, Yale University, New Haven, CT 06520 USA; H. ISHIKAWA, H. TOYOSAKI, and N. MIKAMI, Department of Chemistry, Graduate School of Science, Tohoku University, Aoba-ku, Sendai 980-8578 Japan.

Emission spectra obtained from bulk-gas methinophosphide (HCP) have been interpreted through use of a novel algebraic scheme that explicitly takes into account inherent non-rigidity of the molecular framework.^{a} Fluorescence accompanying selective excitation of individual Ã^{1}A′′ − X̃^{1}Σ^{+} vibronic bands was dispersed under moderate resolution, with the appearance of substantial activity in the ν_{2} bending mode reflecting the bent-from-linear nature of the Ã ← X̃ transition. Aside from providing an economical parameterization for observed patterns of vibrational term energies, the algebraic approach affords a robust and facile means for the quantitative evaluation of multidimensional Franck-Condon factors. These results, as well as subsequent extensions designed to account for non-Condon effects, will be discussed in order to further elucidate the unique structure and dynamics exhibited by participating electronic states.

aH. Ishikawa, H. Toyosaki, N. Mikami, F. Pérez-Bernal, P. H. Vaccaro, and F. Iachello, Chem. Phys. Lett. \underline{365}, 57 (2002).